Shift report 20030319 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Mar 19 2003 - 00:01:50 EST

  • Next message: Peter H. L. Christiansen: "private web pages"
    ------------------------------------------------------------
    Date: Wed Mar 19 00:01:50 2003, Shift: 16:00-24:00
    Supervisor: CEJ, with: 
    Log book#14 pages 68 - 68
    
    Shift Summary:
    ------------------------------------------------------------
    1) Several attemps to store the beams. Succes at 21:30. No 
       beam at BRAHMS before 22:30. High backgrounds (~20) to
       begin with, improved around 23:20 (~6), got bad again  
       around 23:40. 
    
    2) MRS@90(B1050), FS@12(B1/4)
       8494 10K tr 6, 5K tr 3
       8495 58K tr 6, 38K tr 3 (high bg, T3 off)
       8496 still running
    
    3) T3 field, cathode and anode trip. Ended up running with T3
       off. No other problems reported.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Wed Mar 19 2003 - 00:06:17 EST